Solution for x and y
Y=x+4 and
Y=2x^2+6x_8
1
Expert's answer
2020-09-11T09:13:21-0400

"y = x+4\\
y = 2x^2+6x-8\\
x+4 = 2x^2+6x-8 \\
2x^2+6x-8 - x - 4 = 0\\
2x^2+5x-12 = 0\\
x = \frac{-b\pm\sqrt{b^2-4ac}}{2a}\\
x_1 = \frac{-5+\sqrt{5^2-4×2×(-12)}}{2×2}=1.5\\
x_2 = \frac{-5-\sqrt{5^2-4×2×(-12)}}{2×2}=-4\\"


From here, y =:

"y_1 = x_1 + 4 = 1.5 + 4 = 5.5 \\
y_2 = x_2 + 4 = -4 + 4 = 0"


Answer: (1) 1.5 and 5.5, (2) -4 and 0.